The demand and total cost functions of a good are respectively  and

a) Find expressions for TR, (profit) , MR, and MC in terms of Q.

b) Solve the equation

and hence determine the value of Q which maximizes profit.

c) Verify that, at the point of maximum profit, MR=MC.


1
Expert's answer
2021-05-18T19:35:06-0400

a)

Since the two equations for TR and TC are missing for this equation, I will use the following equations to explain this answer.


"TR=-2Q^2+20Q"

"TC=Q^3-8Q^2+20Q+10"

"MR=-4Q+20"

"MC=3Q^2-16Q+20"


We need to set "MR=MC"

"3Q^2-16Q+20=-4Q+20"

"3Q^2-12Q=0."


b)

Solving the equation.

"3Q^2-12Q=0"

"3Q(Q-4)=0"

But we know that "Q>0" for "MC=MR"

Therefore,

"Q-4=0"

"Q=4" .


c)

When the value of the final unit of product (marginal revenue) matches the cost of manufacturing the last unit of product, a manager optimizes profit (marginal cost). As a result, the company will not produce that unit.

Profit is maximized where MR=MC in a perfectly competitive market.

Maximum profit will be expressed as "P=MC" or "P=MR"

"MC=" "-4(4)+20"

"=4"

Maximum profit will be obtained at 4 units.